Hi, Can you provide the exact answers in details for this problem: Calculate the accounting, cash,...

Hi,

Can you provide the exact answers in details for this problem:

Calculate the accounting, cash, and financial break-even quantities.

Here is the problem:

Consider a project to supply Detroit with 27,000 tons of machine screws annually for automobile production. You will need an initial $5,300,000 investment in threading equipment to get the project started; the project will last for 6 years. The accounting department estimates that annual fixed costs will be $1,275,000 and that variable costs should be $240 per ton; accounting will depreciate the initial fixed asset investment straight-line to zero over the 6-year project life. It also estimates a salvage value of $650,000 after dismantling costs. The marketing department estimates that the automakers will let the contract at a selling price of $350 per ton. The engineering department estimates you will need an initial net working capital investment of $510,000. You require a return of 14 percent and face a tax rate of 25 percent on this project. Calculate the accounting, cash, and financial break-even quantities.

Thank you,

Carol
Answer #1
1) Contribution margin per unit = 350-240 = $                 110
Fixed costs = Depreciation+Other fixed operating cost = 5300000/6+1275000 = $     21,58,333
Accounting BEP in units = Total fixed costs/CM unit = 2158333/110 = 19621 Units
2) Cash BEP in units = Cash fixed costs/CM unit = 1275000/110 = 11591 Units
3) Financial BEP in units, is the number of units that will result in 0 NPV.
The PV of known Cash flows are:
INITIAL INVESTMENT:
Cost of the equipment $     53,00,000
Increase in NWC $        5,10,000
Total initial investment $     58,10,000
TERMINAL NON OPERATING CASH INFLOWS:
After tax salvage value = 650000*(1-25%) = $        4,87,500
PV of after tax salvage value = 487500/1.14^6 = $        2,22,098
PV of after tax fixed costs = 1275000*75%*(1.14^6-1)/(0.14*1.14^6) = $     37,18,538
PV OF DEPRECIATION TAX SHIELD:
= (5300000/6)*25%*(1.14^6-1)/(0.14*1.14^6) = $        8,58,747
NPV of known cash flows = -5810000+222098+858747 = $    -84,47,693
The PV of the after tax annual contribution margin should be $     84,47,693
Annual after tax contribution margin = 8447693*0.14*1.14^6/(1.14^6-1) = $     21,72,388
Before tax contribution margin = 1216138/75% = $     28,96,517
Number of units to be sold = 2896517/110 = 26332 Units
CHECK:
Total CM = 26332*110 = $     28,96,517
Depreciation = 5300000/6 = $        8,83,333
Other fixed costs $     12,75,000
NOI $        7,38,183
Tax at 25% $        1,84,546
NOPAT $        5,53,638
Add: Depreciation $        8,83,333
OCF $     14,36,971
PV of OCF = 1436971*(1.14^6-1)/(0.14*1.14^6) = $     55,87,902
PV of after tax salvage value = 487500/1.14^6 = $        2,22,098
Total PV of cash inflows $     58,10,001
Less: Initial investment $     58,10,000
NPV $                      1
Almost 0